OpenStudy (anonymous):

express 5√3x^2 - 2√27x^3 in simplest radical form

OpenStudy (anonymous):

5√3x^3 - 2√27x^3 √27x^3= √3x^3*√9 2√27x^3= 2√3x^3*√9 gives you 5√3x^3 - 2√3x^3*√9 factor out the √3x^3 √3x^3(5 - 2√9) √3x^3(5 - 6) √3x^3(-1) -√3x^3 x^3=x^2*x -x√3x

OpenStudy (anonymous):

omg thank you so much